OpenStudy (anonymous):

Please solve: 15x^2-14x+3

OpenStudy (anonymous):

(3x-1)(5x-3)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

15x^2-14x+3 =15x^2-9x-5x+3 [as 9x+5x=14x] =3x(5x-3)-1(5x-3) =(5x-3)(3x-1)